i need help solving: 3x+2=9
Answer by Theo(13342)

Equation is:

3x + 2 = 9

Subtract 2 from both sides of the equation to get:

3x = 7

Divide both sides of the equation by 3 to get:

x = (7/3)

Confirm by replacing x with (7/3) in your original equation.

3 * (7/3) + 2 = 9

This becomes:

7 + 2 = 9 which becomes:

9 = 9 confirming that the solution is correct.